# CLIC: Contrastive Learning Framework for Unsupervised Image Complexity Representation

### v1

@@ 3.3 Contrastive Learning Framework for IC Representation
We constructed a framework for image complexity representation based on contrastive learning to learn image complexity representations efficiently. We employ two independent encoders: the query encoder and the key encoder. The query encoder is responsible for extracting complex features from inputs. In contrast, the key encoder processes the inputs and generates a 128 -dimension vector, which is used for contrastive learning with the feature output from the query encoder. Following [23], both encoders share the same network structure, and the query encoder performs gradient updating, and the key encoder uses momentum updating:
@@
Where $\theta_{k}$ is the parameter of the key encoder, $\theta_{q}$ is the parameter of the query encoder, and $m\in[0,1)$ is the momentum factor (default 0.999). The purpose of momentum updating is to help the model maintain a stable perception of image complexity during training by keeping the key encoder’s historical information while avoiding losing important information due to training fluctuations. During training, query and key encoders each receive different transformed views as input. Note that the query encoder receives views with a small amount of data augmentation, while the key encoder’s input has additional data augmentation.
In addition, contrastive learning methods typically require a large amount of data to train the model to obtain the desired representation. Therefore, in this work, we sample from several publicly available datasets [25, 10], utilizing the image global entropy, which has the highest correlation with the ground truth, as the sampling basis. The sampling results show that the global entropy of the images in the training set obeys a uniform distribution ($ge\in(0,1)$). Thus, the true image complexity distribution in the training set is also close to a uniform distribution. The image complexity of the training set is close to the uniform distribution, and the model can learn image complexity representations of different difficulties in a balanced way.
@@ 4.1.1 Training Data
Data Collection. Our training images are all derived from a random sampling of Flickr-5B [25] and ImageNe-1M [10], constructing a training set of $\sim$ 1 million images with image entropy obeying a uniform distribution (see Appendix B).
@@ 4.1.2 Implementation Details
Training. The query encoder and key encoder are derived from ImageNet pre-training. The batch size is 128 on 4 GPUs. The initial learning rate is 0.03. A total of 200 epochs are trained, with the learning rate reduced by a factor of 10 at the 120th and 160th epochs. The momentum of updating the key encoder is 0.999. Softmax temperature is 0.07. Stochastic Gradient Descent (SGD) is used to optimize the model, where momentum is set to 0.9 and weight decay is set to 0.0001.
Fine-tuning. We fine-tuning the CLIC query encoder on IC9600 with ICNet [12]. The parameters of the CLIC query encoder are frozen during fine-tuning. The batch size is 128 in 4 GPUs. Stochastic Gradient Descent (SGD) is used to optimize the model. The momentum is set to 0.9, the weight decay is set to 0.0001, and the learning rate is fixed to 0.001. Other settings are consistent with ICNet.
@@ 4.2 Comparison with Previous Results
We compare CLIC with other methods for evaluating image complexity, both unsupervised and supervised, and we use the same metrics as in [12], i.e., PCC and Spearman’s correlation coefficient (SRCC) [5]. The trained CLIC query encoder is used as the backbone of ICNet, and the parameters are frozen and then fine-tuning on IC9600 train set (Tab.2). Note that ICNet employs a two-branch structure, whereas we only use the CLIC query encoder as a separate branch.
@@
unsup. SE 0.534 0.498
NR 0.556 0.541
ED 0.569 0.491
AR 0.571 0.481
UAE 0.651 0.635
super. SAE 0.865 0.860
ComplexityNet† 0.873 0.870
HyperIQA 0.935 0.935
P2P-FM 0.940 0.936
ICNet† 0.947 0.944
ICCORN 0.954 0.951
CLIC Res18 0.890 0.883
Res50 0.893 0.887
Res152 0.899 0.893
Swin-B 0.913 0.908
Table 2: Comparison with previous results. † denotes our reproduction results, and unlabeled results are obtained from the literature [12]. unsup denotes unsupervised learning methods. super denotes supervised learning methods.
The results show that the unsupervised group has the best UAE results, but there is a massive gap between the results and those of the supervised group ICCORN [20]. However, it is worth noting that the unsupervised group methods are traditional hand-designed heuristics. On the other hand, most of the methods in the supervised group are based on deep learning methods and achieve much better performance than the unsupervised methods. This is due to the enormous advantage of deep learning for feature extraction. While CLIC does not perform as well as supervised methods, we aim to address the expensive labeling cost and avoid subjective bias further when we replace the model with a more complex one. Both PCC and SRCC increase to 0.913 and 0.908, which are very close to supervised methods.
@@ 4.3 Ablation of CLIC
Regarding the individual members of CLIC, we set up ablation studies to verify their validity. Each case (a $\sim$ e, CLIC, Tab.3) is initialized with an ImageNet pre-trained model, and after training, the CLIC query encoder is fine-tuning on the IC9600 dataset. The results of the table show the validity of the individual members and demonstrate that there is no conflict between them.
case pre-train IC9600
@@
We sampled untrained 1k images from 10 random classes in the ImageNet train set to do t-SNE visualization (Fig.3). ICNet infers these images to get their IC scores. We extracted the output feature of stage 4 in the CLIC query encoder and downscaled them to 2-dimensional space. We labeled their IC scores (divided into ten bands on $(0,1)$) and class id in ImageNet with color, respectively. We find that CLIC successfully clusters samples with closer ICs without clustering similar samples together. This suggests that CLIC can focus on image complexity representation learning without the interference of attributes such as class. We also visualize the results of case (a) (Tab.3) in Appendix D.1.
@@
We select several images from the MS COCO test set and extract the activation map of their convolution in the last layer of the CLIC query encoder (Fig.4). The results present that CLIC is able to disperse hotspots to most regions of the image in cam rather than focusing on individuals. Such results further illustrate that CLIC is able to focus on the primary attributes of an image rather than individual attributes.
@@ 4.4 Further Analysis
Positive Views Crop Strategy. In Sec.(3.1), we proposed a special strategy for view crop. We believe that for the image complexity representation task, the complexity between pairs of positive samples should be close to each other, so we carefully design three crop types, randomly sampling 1k images from ImageNet and processing them according to these crop schemes, with the crop size decreasing from 240 to 96. We also use the image global entropy to evaluate these views and calculate the PCC between them and the original image entropy (Fig.5). The results show that all three crop strategy have the most appropriate sizes, and we trained CLIC based on these sizes to get the same results as we expected (Tab.4). First, the augmentation methods become more varied, introducing a rich diversity to the training set, and the performance keeps improving. Second, we made two control sets (case b vs. c, case d vs. e) of suitable and unsuitable crop sizes. We found a difference between them, which is small but validates the effectiveness of our proposed crop strategy.
Contributions of Labels. We compare with supervised ICNet by gradually increasing the number of labels used for learning during CLIC fine-tuning, and the results are presented in Fig.6. Note that ICNet is not fine-tuning but trained from scratch with the same number of labels. First, both methods improve their performance as expected with an increasing number of labels. Then, we observe that CLIC has a PCC close to 0.4 from the beginning with only ten labels, while ICNet is only around 0.1. As the number of labels keeps increasing, CLIC gradually approaches saturation at 5k labels since only the header updates the parameters in fine-tuning. In contrast, the parameters in the feature extraction part are already fixed. However, this trend shows that CLIC, after comparative learning with a small number of labels fine-tuning again, can rapidly increase the image complexity characterization results, proving the effectiveness of our method. In the end, we find that the supervised learning method ICNet is much more accurate than CLIC when a sufficiently large number of labels is reached.
case crop&size PCC↑ SRCC↑
@@
Table 4: Comparison of different positive view selections. The crop&size denotes the crop method and the corresponding size. The oc is only crop, fa is few augmentation crop, and ma is more augmentation crop.
@@
Complexity-Aware Loss. To verify the role of global entropy in CAL, we chose three cases of compress ratio, edge density, and None for comparison (Tab.5). We find that the performance without CAL is a relatively low 0.759. After setting the prior information in CAL to edge density and global entropy, the PCC improves by 0.22 and 0.47, respectively, which suggests that metrics with higher relevance to the ground truth as a prior are all effective in improving the performance. Similarly, the low (or even negative) correlation between compress ratio and ground truth leads to training failure. The best performance of global entropy is because the hyper-parameter of edge density is not well chosen. Otherwise, both of them can perform close to each other. In addition, we did an ablation of the $\lambda$ of CAL and obtained the optimal value of 0.25 (Tab.6) .
@@
4.5 Results of Apply to Downstream Tasks.
Object Detection. We apply our CLIC to Mask R-CNN [22] with several backbones. The train details are the same as MMDetection [2]. The results are shown in Table 6. Overall, the addition of image complexity features improves the detection model’s performance.
Backbone       AP       AP50       AP75
R-50 38.2 58.5 41.4
R-10 40.0 60.5 44.0
X-101-32x4d 41.9 62.5 45.9
R-50† 38.5(+0.3) 59.1(+0.6) 41.8(+0.4)
R-101† 40.2(+0.2) 60.9(+0.4) 44.2(+0.2)
X-101-32x4d† 42.0(+0.1) 62.7(+0.2) 45.8(-0.1)
X-101-32x4d†† 42.3(+0.4) 63.3(+0.8) 46.5(+0.6)
Table 7: Comparison of Mask R-CNN on MS COCO. All models are trained with 1× learning rate schedule. † indicate CLIC is applied. The CLIC query encoder weight of feature map which is add to FPN is 0.5. CLIC is a ResNet18. †† means CLIC is a ResNet152.
Method Mem       FPS       mIoU
FCN 5.6G 19.64 66.53
FCN† 14.5G 15.26(-4.38) 68.51 (+1.98)
FCN†† 14.5G 17.83(-1.81) 68.50 (+1.97)
Table 8: Comparison of FCN on PASCAL VOC 2012. The backbone of FCN is R-50-D8. † indicate CLIC is applied. The CLIC query encoder weight of feature map which is add to FCN head is 0.5. CLIC query encoder is a ResNet101. FCN†† indicates the concurrent processing of CLIC and FCN.
Semantic Segmentation. We apply our CLIC to FCN [35] with R-50-D8. The train details are the same as MMSegmentation [8]. Compared to the FCN, the CLIC-assisted FCN† has a mIoU improvement of 1.98. This demonstrates the auxiliary optimization effect of image complexity on semantic segmentation tasks. We visualize some results of the PASCAL VOC 2012 validation set in Fig.7, where the segmentation results of FCN† show a larger and more accurate coverage of the pixels for each object.
Figure 7: Visualization of FCN with PASCAL VOC 2012 val set.
5 Conclusion
Image complexity assessment can aid in optimizing CV tasks. This paper proposes CLIC, an unsupervised image complexity representation learning framework based on comparative learning. Unlike previous approaches that rely on well-labeled datasets, CLIC can efficiently learn IC features based on unlabeled data. Through extensive experimental validation, the results show that CLIC can learn effective IC representations and performs competitively with supervised learning methods for fine-tuning on the IC9600. In addition, CLIC shows significant improvement in several downstream tasks, demonstrating its potential in real-world scenarios. Our future work is to build more suitable pretext tasks for IC representation learning that are entirely free of interference from attributes such as categories, objects, and so on. Furthermore, we also consider developing IC distillation frameworks as a free lunch to optimize downstream tasks.
Supplementary Material
A Preliminary
A.1 Definition of Image Complexity
Image complexity is more difficult to define. Traditional manual methods can be calculated using image entropy, image compression ratio, etc., but most of them are based on heuristic metrics, which are inaccurate to assess and have insufficient generalization ability. Machine learning methods use supervised learning to model image complexity through high annotation costs, a process that inevitably learns human subjective biases. However, whether it is a manual method or a machine learning method, image complexity modeling can be defined as Eq.(7), where ${x_{i}}$ denotes an image from a dataset, and $f(\cdot)$ denotes the mapping from this image to its image complexity $IC_{i}$, with the range of $IC_{i}$ being 0 to 1.
$I{C_{i}}=f({x_{i}})$ (7)
After analyzing the image complexity of several datasets, we observe that, in general, a dataset’s Image Complexity Distribution (ICD) is close to a normal distribution i.e. $I{C_{i}}\sim N(\mu,{\sigma^{2}})$ , indicating that most of the images are concentrated at a medium level of complexity. As shown in Fig.8, the ICD of the dataset is calculated based on image entropy and machine learning methods, respectively, and the distribution curve belongs to the normal distribution category. It is worth mentioning that most of such datasets with normal characteristic ICDs are natural scene images, such as street, indoor and outdoor, and field. However, some datasets, such as the Cornell Grasp Dataset [27], have high frequencies in their ICD concentrated around 0.17, which is one of the reasons why the learning difficulty of these tasks is low (only when Cornell Grasp Dataset is used for 2D target detection task).
Figure 8: Image Complexity Distribution of Datasets. (a) global entropy of MS COCO [32]. (b) ICD of MS COCO [32]. (c) ICD of PASCAL VOC [11]. (d) ICD of Cornell Grasp [27].
A.2 Impaction of Image Complexity
[…]
IC Range AP AP50 AP75
All 40.5 59.3 43.8
<0.3 41.2(+0.7) 60.3(+1.0) 44.6(+0.8)
>0.7 39.6(-0.9) 58.0(-1.3) 42.6(-1.2)
Table 9: Results of YOLOX-S on MS COCO with different IC range. The training details for YOLOX-S were maintained consistent with the official implementation. All represents the original MS COCO val set, with the IC ranger from 0 to 1.
B Datasets
B.1 ImageNet
B.2 Flikr
Flikr [25] contains $\sim$ 5 billion images of multiple types from Flickr (a popular online photo and video hosting, sharing, and management platform), covering a rich diversity of scenes and content. These images range from common scene types, such as natural scenes from everyday life, portraits, buildings, and city streetscapes, to more creatively expressive content, such as still life and artistic photography. The complexity of the images shows significant differences in the number of visual elements, distribution, texture, light and shadow changes, etc., which is suitable for image complexity analysis and evaluation.
B.3 IC9600
C Metrics Details
C.1 Unsupervised Activation Energy
C.2 Edge Density
C.3 Compress Ratio
$CR=\frac{B}{H(I)}\quad\\ \quad H(I)=-\sum_{x=0}^{255}p(x)\log_{2}p(x)$ (10)
D More Ablation
D.1 t-SNE Visualization of Case (a)
@@ D More Ablation
We downscale and visualize the stage 4 output features of CLIC query encoder in Fig.3. The results show that CLIC can learn the image complexity representation well without interference from image categories. Here we create a control group and do the same operation for case (a) in Tab.3 and get the result in Fig.9. It is clear that case (a) is not able to learn the image complexity representation effectively compared to CLIC.
D.2 Correlation of Train Data and Complexity-Aware Loss
@@
We used global entropy as a sampling basis to construct the training set and also used it to design the image complexity-aware loss. Intuitively, it is easy to think there is a strong coupling between them. Therefore, we designed multiple sets of controlled experiments to prove that no such strong coupling relationship exists and that other metrics can achieve closer results (Tab.10). We first build a training set to train CLIC using random sampling, and the results obtained for PCC and SRCC are 0.809 and 0.792. Based on this, we apply CAL, and the PCC and SRCC are increased by 0.032 and 0.035, respectively, which shows that the effect of CAL alone is the enhancement of the results is also very obvious. Further, we also use three metrics, namely, compression ratio, edge density, and global entropy, and we collect the data with uniform and Gaussian distributions. The results show that edge density and global entropy are also close. Conversely, the compress ratio has a negative correlation with ground truth, resulting in poorer results. We sampled with Gaussian distribution because more IC scores in IC9600 are distributed between 0.3 and 0.7. However, it may be because most standard Gaussian distributions are concentrated in 0.4 to 0.6. The number of samples of other complexities is too small, which leads to a decrease in the final results (compared to the uniform distribution).
@@
D.3 Small-scale Crop and Merge
Small-scale Crop and Merge ablation results are presented in Tab.11. We find that as $c$ gets progressively larger, the size of the training set rapidly increases from the original 1× to 18×. Further, we find that the PCC and SRCC reach their highest when $c$ is 3, which are 0.906 and 0.900, respectively. However, the PCC and SRCC begin to decay rapidly as $c$ continues to get larger. The reason for this phenomenon is that as $c$ increases the train set is almost overrun by generated images (which can be considered as pseudo images). And too many pseudo images can seriously interfere with the positive value from real data.
@@ D.3 Small-scale Crop and Merge
D.4 Image Complexity Assessment
We select images (Fig.10) with different difficulty scores and get their image complexity maps and image complexity scores based on pre-trained CLIC and ICNet fine-tuning. The results show that our framework achieves good results in image complexity evaluation at different difficulty levels.
@@ D.4 Image Complexity Assessment
E Future Work

### v2

@@ 4.3 Contrastive Learning Framework for IC Representation
We have developed a framework for image complexity representation based on contrastive learning to efficiently learn image complexity features. The framework employs two independent encoders: the query encoder and the key encoder. The query encoder is tasked with extracting complex features from the input images, while the key encoder processes the inputs and generates a 128-dimensional vector, which is then used for contrastive learning with the feature output from the query encoder. In line with [17], both encoders share the same network architecture, with the query encoder being updated via gradient-based optimization, and the key encoder employing momentum-based updates.
@@
where $\theta_{k}$ is the parameter of the key encoder, $\theta_{q}$ is the parameter of the query encoder, and $m\in[0,1)$ is the momentum factor (default 0.999).
The purpose of momentum updating is to help the model maintain a stable perception of image complexity during training by preserving the key encoder’s historical information, thereby mitigating the risk of losing important details due to fluctuations in the training process. During training, the query and key encoders receive different transformed views as input. Specifically, the query encoder processes views with minimal data augmentation, while the key encoder is provided with inputs that include additional data augmentation.
Furthermore, contrastive learning methods typically require large datasets to train the model and obtain the desired representations. In this work, we sample from several publicly available datasets [43, 44], using the global entropy of the images—which exhibits the highest correlation with the ground truth—as the basis for sampling. The sampling results indicate that the global entropy of the images in the training set follows a uniform distribution ($ge\in(0,1)$), suggesting that the true image complexity distribution in the training set is also close to a uniform distribution. As a result, the model is exposed to a balanced representation of image complexities of varying difficulty levels, facilitating the learning of diverse image complexity representations.
@@ 5.1.1 Training Data
Data Collection. The training images are randomly sampled from Flickr-5B [43] and ImageNet-1M [44], resulting in a training set of approximately 1 million images. The image entropy of these samples follows a uniform distribution.
@@ 5.1.2 Implementation Details
Training. Both the query encoder and the key encoder are pre-trained on ImageNet. We use a batch size of 128 across 4 NVIDIA Geforce 3090 GPUs. The initial learning rate is set to 0.03. Training proceeds for 200 epochs, with the learning rate being reduced by a factor of 10 at the 120th and 160th epochs. The momentum for updating the key encoder is set to 0.999. The softmax temperature is set to 0.07. Stochastic Gradient Descent (SGD) is used for optimization, with momentum set to 0.9 and weight decay set to 0.0001.
Fine-tuning. Fine-tuning of the CLIC query encoder is performed on IC9600 using ICNet [13]. The fine-tuning pipeline is shown in Fig.1 in Appendix C. During fine-tuning, the parameters of the CLIC query encoder are frozen. The batch size is 128 across 4 GPUs. SGD is used for optimization, with momentum set to 0.9, weight decay set to 0.0001, and the learning rate fixed at 0.001. All other settings are consistent with those used in ICNet.
@@ 5.2 Comparison with Previous Results
We compare CLIC with other methods for evaluating image complexity, both unsupervised and supervised, using the same evaluation metrics as in [13], namely the Pearson Correlation Coefficient (PCC) and Spearman’s Rank Correlation Coefficient (SRCC) [45]. The trained CLIC query encoder serves as the backbone of ICNet, with its parameters frozen before fine-tuning on the IC9600 training set (see Tab.3). It is important to note that while ICNet uses a two-branch structure, we only utilize the CLIC query encoder as a single branch.
@@
unsup. SE [46] 0.534 0.498
NR [46] 0.556 0.541
ED [12] 0.569 0.491
AR [47] 0.571 0.481
UAE [29] 0.651 0.635
super. SAE [29] 0.865 0.86
ComplexityNet [15] 0.873 0.870
HyperIQA [14] 0.935 0.935
P2P-FM [48] 0.940 0.936
ICNet [13] 0.947 0.944
ICCORN [16] 0.954 0.951
CLIC Res18 [40] 0.890 0.883
Res50 [40] 0.893 0.887
Res152 [40] 0.899 0.893
Swin-B [41] 0.913 0.908
Table 3: Comparison with previous results. $\dagger$ denotes our reproduction results, and unlabeled results are obtained from the literature [13]. unsup. denotes unsupervised methods. super. denotes supervised methods.
The results indicate that the unsupervised methods achieve the best results in UAE, but there is a significant performance gap when compared to the supervised method, ICCORN [16]. It is important to note that the unsupervised group primarily relies on traditional hand-designed heuristics. In contrast, most methods in the supervised group are based on deep learning, which significantly enhances feature extraction and leads to much better performance. While CLIC does not outperform the supervised methods, its goal is to address the high cost of manual labeling and eliminate subjective bias. When replacing the model with a more complex one, both the PCC and SRCC improve to 0.913 and 0.908, respectively, which are very close to those of the supervised methods.
@@ 5.3 Ablation of CLIC
To verify the validity of the individual components of CLIC, we conducted ablation studies. Each case (from a to e, as shown in Tab.4) begins with an ImageNet pre-trained model. After training, the CLIC query encoder is fine-tuned on the IC9600 dataset. The results presented in the table validate the effectiveness of each individual component and demonstrate that there is no conflict between them.
Case Pretraining IC9600
@@
We randomly sampled 1,000 untrained images from 10 classes in the ImageNet training set to perform t-SNE visualization (Fig.5). ICNet was used to infer the IC scores of these images. We then extracted the output features from stage 4 of the CLIC query encoder and reduced them to a 2-dimensional space. The images were labeled by their IC scores (divided into ten bands in the range $(0,1)$) and ImageNet class IDs, with distinct colors for each. The results show that CLIC effectively clusters samples with similar ICs, without grouping images from the same class together. This indicates that CLIC can focus on learning image complexity representations without being influenced by attributes like class. Additionally, we visualize the results of case (a) (Tab.4) in Fig.9.
@@
We selected several images from the MS COCO test set and extracted the activation maps of their convolutions from the last layer of the CLIC query encoder (Fig.6). The results show that CLIC disperses hotspots across most regions of the image in the class activation map (CAM), rather than concentrating on specific objects. This demonstrates that CLIC is capable of focusing on the primary attributes of an image, rather than being biased toward individual attributes.
@@ 5.4 Further Analysis
Positive Views Crop Strategy. In Sec.4.1, we introduced a special strategy for cropping views. For the image complexity representation task, we believe that the complexity between pairs of positive samples should be closely matched. To achieve this, we carefully designed three crop types by randomly sampling 1k images from ImageNet and processing them with crop sizes ranging from 240 to 96. We also utilized the image global entropy to evaluate these views and computed the Pearson correlation coefficient (PCC) between them and the original image entropy (Fig.7). The results indicate that all three crop strategies were appropriately sized, and when CLIC was trained based on these sizes, we obtained the expected results (Tab.5). First, the augmentation methods introduced more variety, enriching the diversity of the training set, leading to continuous performance improvements. Second, we created two control sets (case b vs. c, case d vs. e) using suitable and unsuitable crop sizes. Although the difference between them was small, it effectively validated the usefulness of our proposed crop strategy.
Contributions of Labels. We compared CLIC with supervised ICNet by gradually increasing the number of labels used during CLIC fine-tuning, with the results presented in Fig.8. Note that ICNet was trained from scratch with the same number of labels, rather than fine-tuned. Initially, both methods show improved performance with an increasing number of labels. Notably, CLIC achieves a PCC close to 0.4 with just ten labels, while ICNet only reaches around 0.1. As the number of labels increases, CLIC gradually approaches saturation after 5k labels, since only the header is updated during fine-tuning, while the feature extraction parameters remain fixed. However, this trend demonstrates that CLIC can rapidly improve image complexity characterization with even a small number of labeled samples, validating the effectiveness of our approach. Ultimately, we observe that when a sufficiently large number of labels is used, the supervised learning method (ICNet) significantly outperforms CLIC in accuracy.
case aug & crop PCC↑ SRCC↑
@@
Table 5: Comparison of different positive view selections. The aug & crop denotes the augmentation type and crop type with the corresponding size. The oc is only crop, fa is few augmentation crop, and ma is more augmentation crop.
@@
Complexity-Aware Loss. To evaluate the role of global entropy in CAL, we compared three cases: compress ratio, edge density, and no prior information (Tab.6). The results show that the performance without CAL is relatively low, with a PCC of 0.759. When edge density and global entropy were used as prior information in CAL, the PCC improved by 0.22 and 0.47, respectively, indicating that metrics with higher relevance to the ground truth are effective in enhancing performance. Conversely, the compress ratio, which has a low (or even negative) correlation with the ground truth, led to training failure. The best performance was achieved with global entropy, as the hyperparameter for edge density was not well optimized. If the edge density hyperparameter were chosen better, its performance would be closer to global entropy. Additionally, we conducted an ablation study on the $\lambda$ parameter of CAL and identified the optimal value as 0.25 (Tab.7).
@@
t-SNE Visualization of Case (a). We downscale and visualize the stage 4 output features of the CLIC query encoder, as shown in Fig.5. The results indicate that CLIC effectively learns the image complexity representation without interference from image categories. For comparison, we create a control group and perform the same procedure for case (a) in Tab.5, with the result presented in Fig.9. It is evident that case (a) fails to learn the image complexity representation effectively when compared to CLIC.
Figure 9: t-SNE visualization of case (a). All feature are taken from the output of stage 4 in CLIC query encoder (ResNet). The complexity score (a) is divided into ten segments. The class id (b) is from 0 to 9.
@@
Decoupling Analysis of Training Data and CAL. Global entropy was used as both a sampling basis for constructing the training set and as part of the image complexity-aware loss function. Intuitively, one might expect a strong coupling between these two components. To investigate this, we designed multiple controlled experiments to demonstrate that no such strong coupling exists and that other metrics can also yield comparable results (Tab.8). Initially, we built a training set using random sampling for CLIC, and the resulting PCC and SRCC were 0.809 and 0.792, respectively. When applying CAL, the PCC and SRCC increased by 0.032 and 0.035, respectively, indicating that CAL alone significantly enhances the results. Additionally, we tested three metrics—compression ratio, edge density, and global entropy—by sampling data from both uniform and Gaussian distributions. The results showed that edge density and global entropy were similar in their performance. In contrast, the compression ratio had a negative correlation with the ground truth, leading to poorer results. We sampled using a Gaussian distribution because most IC scores in the IC9600 dataset are concentrated between 0.3 and 0.7. However, the standard Gaussian distribution primarily concentrates between 0.4 and 0.6, which resulted in fewer samples of other complexities and consequently reduced the final results compared to the uniform distribution.
@@
Small-scale Crop and Merge. The results are shown in Tab.9. As the value of $c$ increases, the size of the training set expands rapidly, growing from the original 1× to 18×. We observe that the highest PCC and SRCC are achieved when $c$ is set to 3, with values of 0.906 and 0.900, respectively. However, both PCC and SRCC begin to decline significantly as $c$ increases further. This phenomenon can be explained by the fact that as $c$ grows, the training set becomes dominated by generated images, which can be considered pseudo images. The overabundance of pseudo images interferes with the contribution of real data, thereby reducing the model’s performance.
@@
Image Complexity Assessment. We select images with varying difficulty scores (Fig.10) and compute their image complexity maps and scores using the pre-trained CLIC model and ICNet after fine-tuning. The results demonstrate that our framework performs effectively in evaluating image complexity across different levels of difficulty.
5.5 Results of Apply to Downstream Tasks.
Application Pipelines. The pipelines for applying to downstream tasks are shown in Fig.11. Mask R-CNN and FCN are used as baseline models for the detection and segmentation tasks, respectively.
Figure 11: Visualization of FCN with PASCAL VOC 2012 val set.
Object Detection. We apply CLIC to Mask R-CNN [49] with several backbones. The training details follow those in MMDetection [50]. The results are shown in Table 10. Overall, incorporating image complexity features improves the performance of the detection model. However, as the backbone of Mask R-CNN becomes deeper, the performance improvement brought by CLIC diminishes. We hypothesize that this is due to the ResNet18-based CLIC network having fewer layers and thus being less capable of extracting image features compared to ResNet50 and ResNet101. To address this, we apply CLIC based on ResNet152 to the X-101-32x4d [51] (Tab.10 X-101-32x4d††). We expect that a deeper network will better capture image complexity features. The results confirm our hypothesis, with the performance improvements for X-101-32x4d†† surpassing those of X-101-32x4d.
Backbone      AP      AP50      AP75
R-50 [40] 38.2 58.5 41.4
R-10 [40] 40.0 60.5 44.0
X-101-32x4d [51] 41.9 62.5 45.9
R-50 [40] 38.5(+0.3) 59.1(+0.6) 41.8(+0.4)
R-101 [40] 40.2(+0.2) 60.9(+0.4) 44.2(+0.2)
X-101-32x4d [51] 42.0(+0.1) 62.7(+0.2) 45.8(-0.1)
X-101-32x4d†† [51] 42.3(+0.4) 63.3(+0.8) 46.5(+0.6)
[…]
Method Mem       FPS       mIoU
FCN [52] 5.6G 19.64 66.53
FCN [52] 14.5G 15.26(-4.38) 68.51 (+1.98)
FCN†† [52] 14.5G 17.83(-1.81) 68.50 (+1.97)
Table 11: Comparison of FCN on PASCAL VOC 2012. The backbone of FCN is R-50-D8. $\dagger$ indicate CLIC is applied. The CLIC query encoder weight of feature map which is add to FCN head is 0.5. CLIC query encoder is a ResNet101. FCN†† indicates the concurrent processing of CLIC and FCN.
Figure 12: Visualization of FCN with PASCAL VOC 2012 val set.
Semantic Segmentation. We apply CLIC to FCN [52] with R-50-D8. The training details follow those in MMSegmentation [53]. Compared to the baseline FCN, the CLIC-assisted FCN shows an improvement of 1.98 mIoU, demonstrating the beneficial effect of image complexity optimization on semantic segmentation tasks. However, we observe that CLIC increases GPU memory consumption by approximately 9GB, and the FPS drops by 4.38. This represents a drawback of CLIC. Both training and inference in FCN require the use of CLIC, as the knowledge from CLIC cannot be transferred directly to FCN. While concurrent execution (FCN††) can improve the running speed, the improvement is minimal. Additionally, while GPU memory usage is similar for both concurrent and sequential execution, the CPU and regular memory must handle the extra load, which is undesirable for practical engineering applications. This is an area we plan to address in future work. We visualize some results from the PASCAL VOC 2012 validation set in Fig.12, where the segmentation results from FCN show a larger and more accurate coverage of pixels for each object.
6 Limitations